# A multiple imputation approach to distinguish curative from life-prolonging effects in the presence of missing covariates

### v1

@@ Abstract
Medical advancements have increased cancer survival rates and the possibility of finding a cure. Hence, it is crucial to evaluate the impact of treatments in terms of both curing the disease and prolonging survival. We may use a Cox proportional hazards (PH) cure model to achieve this. However, a significant challenge in applying such a model is the potential presence of partially observed covariates in the data. We aim to refine the methods for imputing partially observed covariates based on multiple imputation and fully conditional specification (FCS) approaches. To be more specific, we consider a more general case, where different covariate vectors are used to model the cure probability and the survival of patients who are not cured. We also propose an approximation of the exact conditional distribution using a regression approach, which helps draw imputed values at a lower computational cost. To assess its effectiveness, we compare the proposed approach with a complete case analysis and an analysis without any missing covariates. We discuss the application of these techniques to a real-world dataset from the BO06 clinical trial on osteosarcoma.
@@ 1 Introduction
Recent advancements in medicine have increased the chances of survival for some types of cancer such as breast cancer, melanoma, osteosarcoma, and childhood acute lymphocytic leukemia, and have even made cure possible. As a result, it is important not only to estimate the impact several risk factors may have on survival but also to analyze their effect on the probability of being cured. This is especially relevant in pediatric oncology. Osteosarcoma, a type of bone tumour mainly affecting children and young adults, may help illustrate this need. For this type of cancer, over 95% of patients who do not experience a relapse within 5 years after treatment are considered cured and are unlikely to relapse later (Ferrari et al., 2006). Hence, it is crucial to consider cured patients in statistical analyses and to assess the efficacy of a treatment in terms of cure rather than only focusing on its impact on prolonging survival.
@@
In survival analysis, individuals who will never experience the event of interest, such as disease recurrence or disease-related death, are referred to as "cured". Traditional survival analysis methods can be less informative when a portion of the population is considered cured, as they may not account for differences in cure status within the population. Additionally, the cure status is not always observable; while we know that patients who experience the event of interest are uncured, we cannot determine if censored patients will eventually experience the event. This led to the development of cure models, which provide a more suitable modelling approach, particularly in oncological studies (Legrand and Bertrand, 2019).
@@
One type of cure model is the mixture cure model, which defines overall survival as a combination of the cure probability (incidence) and the survival for uncured patients (latency). Mixture cure models allow for the use of different covariates when modelling the incidence and the latency components. This means we can differentiate between the effect of a covariate on the probability of being cured and its effect on the survival of uncured patients. In this context, we consider a Cox proportional hazards (PH) cure model, where the incidence is modelled by a logistic regression model and a Cox PH model is employed for the latency. Estimation is generally carried out using a maximum likelihood approach in an incomplete data setting. For this purpose, the expectation-maximization (EM) algorithm is often used due to the latent cure status (Cai et al., 2012). For more detailed information on cure models, see the review paper by Amico and Van Keilegom (Amico and Van Keilegom, 2018) or the book of Peng and Yu (Peng and Yu, 2021).
@@
A recent study by Musta et al. (Musta et al., 2022) examined the prognostic significance of histologic response and intensified chemotherapy on the cure rate and progression-free survival among osteosarcoma patients who have not been cured. However, the study excluded observations with missing histological responses from the analysis. In this article, we propose a method based on a multiple imputation approach, to analyze a "completed" dataset.
@@
When working with an incomplete dataset, a simple but naive approach to handling missing data is to conduct a complete-case analysis, which involves removing observations with missing data. However, this method can lead to inefficient and inconsistent results when the missing mechanism is not completely at random (Little and Rubin, 2002). Alternatively, missing values can be filled in by using simple imputation techniques, either based on summaries (e.g. mean) or statistical models (e.g. regression) based on available data. While these methods allow for the use of the whole sample, completed by imputation, they may underestimate the final model’s standard errors (van Buuren, 2018). To address these issues, Rubin (Rubin, 1987) introduced multiple imputation techniques. This approach produces unbiased estimates and incorporates the uncertainty of imputation into standard error estimates. Multiple imputation involves creating multiple imputed datasets, estimating parameters in each of them separately, and then combining the results. When dealing with survival data, it has been found beneficial to include the survival outcome in the imputation model for the covariates (White and Royston, 2009).
@@
However, when dealing with cured patients, the latent cure status presents an additional challenge for the imputation procedure. Beesley et al. (Beesley et al., 2016) have developed a method for using multiple imputation in the Cox proportional hazards (PH), which is based on specifying an exact or approximated distribution of the missing covariates, conditional on the available ones and the survival response (time, status). However, this method assumes that the same design vector is used in the models for both the incidence and latency of the cure, which can be a limitation as including all available covariates in both components of the mixture model may lead to the estimation of a large number of parameters. Additionally, the factors affecting the probability of being cured may differ from those influencing the hazard, making the settings of incidence and latency distinct from each other. Lastly, as demonstrated by Hanin and Huang (Hanina and Huang, 2014), sharing covariates between various components of a Cox PH cure model may prevent the model identifiability. Hence, a simpler model would be preferred (over the one with all the covariates in both components) whenever supported by the data.
In this paper, we propose to release this constraint and to generalize the existing methodology for multiple imputation in cure models by accommodating possibly different sets of covariates when modelling the cure probability and the survival of the uncured patients. This allows us to fully leverage the benefits of a mixture model. Our motivation for this improvement also stems from its application to the analysis of osteosarcoma data (Musta et al., 2022). The results of our simulation study also emphasise the need to differentiate between the covariates modelling the cure probability and the survival of the uncured patients. Particularly, we observe that, when a covariate with missing values is wrongly included in both models, multiple imputation based on the approximate method performs better than the exact conditional distribution approach.
@@
In Section 2, we introduce the fundamental concepts underlying cure models and the use of multiple imputation in these models. In Section 3, we outline the methodology for performing multiple imputation in cure models when the sets of covariates for incidence and latency models are different. In Section 4, we evaluate the performance of this method using simulated data. In Section 5, we apply the developed methodology to analyze osteosarcoma data from the BO06 clinical trial and compare the results with the complete-case approach for sensitivity. Finally, in Section 6, we conclude with a discussion.
@@ 2.1 Mixture cure models
We are interested in studying the time $T$ until an event of interest occurs, in a sample consisting of both cured and uncured individuals. The cured individuals will not experience the event of interest during their lifetime. We denote by $G$ the binary indicator for the uncured status, i.e. $G=1$ if uncured and $G=0$ if cured. However, due to censoring at time $C$ and to the limited time window covered by the study at hand, these variables can not be directly observed. Instead, we observe the follow-up time $Y=\min(T,C)$ and the event indicator $\Delta=I(T<C)$. For uncensored observations ($\Delta=1$), $G=0$, while for censored observations $G$ is unknown.
@@
The survival of the uncured patients, known as latency, given a set of covariates $Z\in\mathbb{R}^{q}$, which may differ from $X$, is represented by $S_{u}(t|Z)=P(T>t|Z,G=1)$, and it can be modelled employing a regression model such as the Cox PH model (Cox, 1972)
@@
where $H_{0}$ denotes the cumulative baseline hazard function of the uncured and $\beta\in\mathbb{R}^{q}$ is the vector of regression parameters. In this model, the proportional hazards assumption holds for the sub-population of uncured patients, but not for the entire population. Combining these components, see (Kuk and Chen, 1992), the overall survival is given by
@@
For a cure model to be identifiable, a sufficiently long follow-up is required, i.e. the study duration is longer than the support of the distribution for the event times of the uncured individuals. For a more extensive discussion regarding the assumption of sufficient follow-up, we refer the reader to the review by Amico and Van Keilegom (Amico and Van Keilegom, 2018). In practice, to check whether such an assumption is satisfied, one typically combines medical knowledge with a visual inspection of the Kaplan-Meier estimator. Ideally, we would like to observe a long plateau after the last observed event, containing many censored observations. Recently, some statistical tests have also been proposed for the sufficient follow-up assumption (Xie et al., 2023; Yuen and Musta, 2024; Maller et al., 2024).
@@
$\mathcal{L}_{c}(\alpha_{0},\alpha,\beta,H_{0})=\prod_{i=1}^{n}[\pi(X_{i})f_{u}(Y_{i}|Z_{i})]^{\Delta_{i}G_{i}}[\pi(X_{i})S_{u}(Y_{i}|Z_{i})]^{(1-\Delta_{i})G_{i}}[1-\pi(X_{i})]^{(1-\Delta_{i})(1-G_{i})}$ (4)
@@
Since this model depends on the latent cure status $G$, the set of parameters is usually estimated using an Expectation-Maximization (EM) algorithm (Peng and Dear, 2000; Sy and Taylor, 2000), which is an iterative two-step procedure. In the first step, the posterior expectation of the complete data log-likelihood over the latent uncured indicator is obtained, conditional on the parameters estimated from the previous iteration. For individual $i\in(1,...n)$ at iteration $r$, this is given by
$q_{i}^{(r)}$ $=E(G_{i}|Y_{i},\Delta_{i},X_{i},Z_{i},\alpha_{0}^{(r-1)},\alpha^{(r-1)},\beta^{(r-1)},H_{0}^{(r-1)})$
$=\Delta_{i}+(1-\Delta_{i})\frac{\pi^{(r-1)}(X_{i})S_{u}^{(r-1)}(Y_{i}|Z_{i})}{1-\pi^{(r-1)}(X_{i})+\pi^{(r-1)}(X_{i})S_{u}^{(r-1)}(Y_{i}|Z_{i})}$
where $\pi^{(r-1)}$ and $S_{u}^{(r-1)}$ are computed according to the equations (1) and (2) by using the parameter estimates resulting from iteration $r-1$, namely $\alpha_{0}^{(r-1)},\alpha^{(r-1)},\beta^{(r-1)},H_{0}^{(r-1)}$. The second step of the EM algorithm consists of maximizing the expected log-likelihood for complete data given in (4), where $G_{i}$ is replaced by its expectation $q_{i}^{(r)}$ in each iteration. This reduces to a simpler weighted data problem, with an expected log-likelihood which can be maximized by standard algorithms.
If parameter distinctiveness holds, the two components of the expected likelihood can be maximized separately, one with respect to $\alpha_{0}$ and $\alpha$, and the other with respect to $\beta$ and $H_{0}$. Maximization with respect to $H_{0}$ at iteration $r$ leads to the following modified Breslow estimator (Breslow, 1974) of the baseline cumulative hazard function
@@ 2.2 Handling missing values by multiple imputation
We first illustrate the multiple imputation approach for a general model with outcome $Y$ and design vector $X\in\mathbb{R}^{p}$, where $X^{(j)}$ has missing observations. The procedure consists of imputing the incomplete covariate $X^{(j)}$ defining a model for this element, conditional on the outcome $Y$ and all the remaining covariates $X^{(-j)}=X\setminus X^{(j)}$. This is done through the fully conditional specification (FCS) (van Buuren et al., 2006) approach, requiring the specification of the conditional distribution $f(X^{(j)}|Y,X^{(-j)};\Psi^{j}$), where $\Psi^{j}$ represents a set of parameters. The procedure begins by fitting a regression model for $X^{(j)}$ with covariates $Y$ and $X^{(-j)}$ on the portion of observed data for $X^{(j)}$. This provides us with estimates $\hat{\Psi}^{(j)}$ for the regression coefficients $\Psi\in\mathbb{R}^{p}$, as well as the estimated covariance matrix $\hat{\Sigma}_{\Psi}$. We then draw values $\Psi^{*}$ from the posterior distribution of $\Psi$ given the observations, which is frequently assumed to be $N(\hat{\Psi}^{(j)},\hat{\Sigma}_{\Psi})$. Imputed values for the cases where $X^{(j)}$ is missing are drawn using this conditional distribution $f(X^{(j)}|Y,X^{(-j)};\hat{\Psi}^{*})$. This step is repeated $K$ times, resulting in $K$ imputed samples.
When there are multiple variables with missing observations, this process is repeated sequentially for each one of them, using the observed and the imputed values of the remaining variables. This procedure, known as multiple imputation by chained equations, is summarized in Figure 1. It starts by filling all missing values randomly and it then creates $K$ datasets with imputed values for the incomplete variables. The order in which the variables are imputed is determined by the proportion of missing cases, starting with the variable that has the least missing data. (van Buuren, 2018). Each variable uses the observed and the most recently imputed values of the remaining ones, and this method is iterated on each dataset until convergence.
@@
After obtaining $K$ imputed datasets, the set of parameters $\Psi$ is estimated separately using each of these datasets, resulting in a vector of estimates $(\hat{\Psi}_{1},...,\hat{\Psi}_{K})$ and estimated covariance matrices $(\hat{\Sigma}_{1},...,\hat{\Sigma}_{K})$. Finally, the pooled estimate is obtained by combining the $K$ estimates using Rubin’s rules (Rubin, 2004)
@@
$\hat{\Sigma}_{\Psi}=\frac{1}{K}\sum\limits_{k=1}^{K}\Sigma_{k}+\left(1+\frac{1}{K}\right)\frac{1}{K-1}\sum\limits_{k=1}^{K}(\Psi_{k}-\hat{\Psi})(\Psi_{k}-\hat{\Psi})^{T}$

### v2

@@ Abstract
Medical advances have increased cancer survival rates and the possibility of finding a cure. Hence, it is crucial to evaluate the impact of treatments both in terms of cure and prolongation of survival. To achieve this, we may use a Cox proportional hazards (PH) cure model. However, a significant challenge in applying such a model is the potential presence of partially observed covariates. We aim to refine the methods for imputing partially observed covariates based on multiple imputation and fully conditional specification (FCS) approaches. To be more specific, we consider a general case in which different covariate vectors are used to model the probability of cure and the survival of patients who are not cured. We investigated the performance of the multiple imputation procedure based on the exact conditional distribution and an approximate imputation model, which helps to draw imputed values at a lower computational cost. To assess the effectiveness of these approaches, we compare them with a complete case analysis and an analysis that includes all available covariates in modelling both cure probabilities and the survival of the uncured. We discuss the application of these techniques to a real-world dataset from the BO06 clinical trial on osteosarcoma.
@@ 1 Introduction
Recent advances in medicine have increased the chances of survival for some types of cancer, such as breast cancer, melanoma, osteosarcoma, and acute childhood lymphocytic leukemia, and have even made cure possible. As a result, it is important not only to estimate the impact that several risk factors may have on survival, but also to analyse their effect on the probability of being cured. This is especially relevant in paediatric cancer. Osteosarcoma, a type of bone tumour that primarily affects children and young adults, may help illustrate this need. For this type of cancer, more than 95% of patients who do not experience a relapse within 5 years after treatment are considered cured and are unlikely to relapse later (Ferrari et al., 2006). In survival analysis, individuals who will never experience the event of interest, such as disease recurrence or disease-related death, are referred to as "cured". Hence, it is crucial to consider cured patients in statistical analyses to assess the efficacy of a treatment in terms of cure rather than only focussing on its impact on prolonging survival.
@@
Traditional survival analysis methods can be less informative when a portion of the population is considered cured, as they may not account for differences in cure status within the population. Furthermore, the cure status is not always observable; while we know that patients who experience the event of interest are uncured, we cannot determine whether censored patients will eventually experience the event. This led to the development of cure models, which provide a more suitable modelling approach, particularly in oncological studies (Legrand and Bertrand, 2019).
@@
One type of cure model is the mixture cure model, which defines overall survival as a combination of the probability of cure (incidence) and the survival of uncured patients (latency). Mixture cure models allow for the use of different covariates when modelling the incidence and latency components. This means that we can differentiate between the effect of a covariate on the probability of being cured and its effect on the survival of uncured patients. In this context, we consider a Cox proportional hazards (PH) cure model, where the incidence is modelled by a logistic regression model and a Cox PH model is used for the latency. Estimation is generally performed using the maximum likelihood approach in an incomplete data setting. For this purpose, the expectation-maximization (EM) algorithm is often used due to the latent cure status (Cai et al., 2012). For more detailed information on cure models, see the review paper by Amico and Van Keilegom (2018) or the book by Peng and Yu (2021).
@@
A recent study by Musta et al. (2022) examined the prognostic importance of the histological response and intensified chemotherapy on the cure rate and progression-free survival in a sample of osteosarcoma patients who have not been cured. However, the study excluded observations with missing histological responses from the analysis. This motivated us to investigate, in this article, the sensitivity of the results to the method used to handle missing values in the histological response.
@@
When working with an incomplete dataset, a simple but naive approach to handling missing data is to conduct a complete case analysis, which involves removing observations with missing data. However, this method can lead to inefficient and inconsistent results when the missing mechanism is not completely at random (Little and Rubin, 2002). Alternatively, missing values can be filled in by using simple imputation techniques, either based on summaries (e.g. mean) or statistical models (e.g. regression) based on available data. While these methods allow for the use of the whole sample, completed by imputation, they may underestimate the final model’s standard errors (van Buuren, 2018). To address these issues, Rubin (1987) introduced multiple imputation techniques. This approach produces unbiased estimates and incorporates the uncertainty of imputation into standard error estimates. Multiple imputation involves creating multiple imputed datasets, estimating the parameters in each of them separately, and then combining the results. When handling survival data, it has been found beneficial to include the survival outcome in the imputation model for the covariates (White and Royston, 2009).
@@
However, dealing with cured patients presents an additional challenge for the imputation procedure. Beesley et al. (2016) have developed a method for using multiple imputation in the Cox proportional hazards (PH) cure model, which is based on specifying either an exact or an approximate distribution of the missing covariates, conditional on the available ones and the survival response (time, status). The authors state in that paper that, for simplicity, they assume the same design vector is used in the models for both incidence and latency. This assumption restricts the applicability of the method in practice, as including all available covariates in both components of the mixture model may lead to the estimation of a large number of parameters. In addition, the factors affecting the probability of being cured can differ from those influencing the hazard, making the incidence and latency settings distinct from each other. Moreover, as demonstrated by Hanin and Huang (2014), the sharing of covariates between various components of a Cox PH cure model may prevent the model identifiability. Hence, a simpler model could be preferred (over the one with all covariates in both components) whenever supported by the data.
In this paper, we consider the more general case of a mixture cure model with possibly different sets of covariates for the incidence and latency component. This allows us to fully leverage the benefits of a mixture model. Specifically, we consider three cases depending on whether the variable with missing values affects only the probability of cure, only the survival of the uncured, or both components. Building on the work of Beesley et al. (2016), we derive approximated imputation models for these scenarios. In particular, only for the latter one, the approximate model is the same as in Beesley et al. (2016). Through simulation studies, we investigate how these two multiple imputation approaches perform in practice in this more general setting, an aspect not explored in the original work by Beesley et al. (2016). The simulation results also emphasise the need to differentiate between the covariates modelling the probability of cure and the survival of the uncured patients, as including all covariates in both components leads to wider confidence intervals. However, model selection is not a trivial task in the presence of missing values and model misspecification seems to deteriorate the performance of the exact imputation method. However, the approximate imputation model appears to be quite robust to model misspecification and would be the recommended choice despite the slightly wider confidence intervals compared to the exact approach. Furthermore, to facilitate the application of our method in practice, we provide general-purpose software [^1] The source code is available at https://github.com/martacip/mi_curemodels.
@@
In Section 2, we introduce the fundamental concepts underlying cure models and the use of multiple imputation in this context. In Section 3, we outline the methodology for performing multiple imputation in cure models when the sets of covariates for incidence and latency models are different. In Section 4, we evaluate the performance of this method using simulated data. In Section 5, we apply the developed methodology to analyse osteosarcoma data from the BO06 clinical trial and compare the results with the complete case approach for sensitivity. Finally, in Section 6, we conclude with a discussion.
@@ 2.1 Mixture cure models
We are interested in studying the time $T$ until an event of interest occurs, in a sample consisting of cured and uncured individuals. The cured individuals will not experience the event of interest during their lifetime. We denote by $G$ the binary indicator of the uncured status, i.e. $G=1$ if uncured and $G=0$ if cured. However, due to the censoring at time $C$ and to the limited time window covered by the study at hand, these variables cannot be directly observed. Instead, we observe the follow-up time $Y=\min(T,C)$ and the event indicator $\Delta=I(T<C)$. For uncensored observations ($\Delta=1$), $G=1$, while for censored observations $G$ is unknown.
@@
The survival of the uncured patients, known as latency, conditional on a set of covariates $Z\in\mathbb{R}^{q}$, which may differ from $X$, is represented by $S_{u}(t|Z)=P(T>t|Z,G=1)$, and it can be modelled employing a regression model such as the Cox PH model (Cox, 1972)
@@
where $H_{0}$ denotes the cumulative baseline hazard function of the uncured and $\beta\in\mathbb{R}^{q}$ is the vector of regression parameters. In this model, the proportional hazards assumption holds for the subpopulation of uncured patients, but not for the entire population. Combining these components (Kuk and Chen, 1992), the overall survival is given by
@@
For a cure model to be identifiable, a sufficiently long follow-up is required, i.e. the study duration should be longer than the support of the distribution for the event times of the uncured individuals. For a more extensive discussion of the assumption of sufficient follow-up, we refer the reader to the review by Amico and Van Keilegom (2018). In practice, to check whether such an assumption is satisfied, one typically combines medical knowledge with a visual inspection of the Kaplan-Meier estimate. Ideally, we would like to observe a long plateau after the last observed event, which contains many censored observations. Recently, some statistical tests have also been proposed for the sufficient follow-up assumption (Xie et al., 2023; Yuen and Musta, 2024; Maller et al., 2024).
@@
$\mathcal{L}_{c}(\alpha_{0},\alpha,\beta,H_{0})=\prod_{i=1}^{n}\left[\pi(X_{i})f_{u}(Y_{i}|Z_{i})\right]^{\Delta_{i}G_{i}}\left[\pi(X_{i})S_{u}(Y_{i}|Z_{i})\right]^{(1-\Delta_{i})G_{i}}\left[1-\pi(X_{i})\right]^{(1-\Delta_{i})(1-G_{i})}$ (4)
@@
Since this model depends on the latent cure status $G$, the set of parameters is usually estimated using an Expectation-Maximization (EM) algorithm (Peng and Dear, 2000; Sy and Taylor, 2000), which is an iterative two-step procedure. In the E-step, the posterior expectation of the complete data log-likelihood over the latent uncured indicator is obtained, conditional on the parameters estimated from the previous iteration. For individual $i\in(1,...n)$ at iteration $r$, this is given by
$\begin{split}q_{i}^{(r)}&=E(G_{i}|Y_{i},\Delta_{i},X_{i},Z_{i},\alpha_{0}^{(r-1)},\alpha^{(r-1)},\beta^{(r-1)},H_{0}^{(r-1)})\\ &=\Delta_{i}+(1-\Delta_{i})\frac{\pi^{(r-1)}(X_{i})S_{u}^{(r-1)}(Y_{i}|Z_{i})}{1-\pi^{(r-1)}(X_{i})+\pi^{(r-1)}(X_{i})S_{u}^{(r-1)}(Y_{i}|Z_{i})}\end{split}$
where $\pi^{(r-1)}$ and $S_{u}^{(r-1)}$ are computed according to equations (1) and (2) by using the parameter estimates resulting from iteration $r-1$, namely $\alpha_{0}^{(r-1)},\alpha^{(r-1)},\beta^{(r-1)},H_{0}^{(r-1)}$. The M-step of the EM algorithm consists of maximising the expected log-likelihood for complete data given in (4), where $G_{i}$ is replaced by its expectation $q_{i}^{(r)}$ at each iteration. This reduces to a simpler weighted data problem, with an expected log-likelihood which can be maximised by standard algorithms.
If parameter distinctiveness holds, the two components of the expected likelihood can be maximised separately, one with respect to $\alpha_{0}$ and $\alpha$, and the other with respect to $\beta$ and $H_{0}$. Maximisation with respect to $H_{0}$ at iteration $r$ leads to the following modified Breslow estimator (Breslow, 1974) of the baseline cumulative hazard function
@@ 2.2 Handling missing values by multiple imputation
We first illustrate the multiple imputation approach for a general model with outcome $Y$ and design vector $X\in\mathbb{R}^{p}$, where $X^{(j)}$ has missing observations. The procedure consists of imputing the incomplete covariate $X^{(j)}$ by defining a model for this element, conditional on the outcome $Y$ and the remaining covariates $X^{(-j)}=X\setminus X^{(j)}$. This is done using the fully conditional specification (FCS) approach (van Buuren et al., 2006), which requires the specification of the conditional distribution $f(X^{(j)}|Y,X^{(-j)};\theta$), where $\theta$ represents a set of parameters. Such parameters depend on $j$, but we will omit it from the notation for simplicity. The procedure begins by fitting a regression model for $X^{(j)}$ with covariates $Y$ and $X^{(-j)}$ on the portion of observed data for $X^{(j)}$. This provides us with estimates $\hat{\theta}$ for the regression coefficients $\theta\in\mathbb{R}^{p}$, as well as the estimated covariance matrix $\hat{\Sigma}_{\theta}$. We then draw values $\theta^{*}$ from the posterior distribution of $\theta$ given the observations, which is approximated by a Gaussian distribution, $N(\hat{\theta},\hat{\Sigma}_{\theta})$. The imputed values for the cases where $X^{(j)}$ is missing are drawn from the conditional distribution $f(X^{(j)}|Y,X^{(-j)};\hat{\theta}^{*})$. This step is repeated $K$ times, resulting in $K$ imputed samples.
When there are multiple variables with missing observations, this process is repeated sequentially for each one of them, using the observed and imputed values of the remaining variables. This procedure, known as multiple imputation by chained equations, is summarised in Figure 1. It starts by filling all missing values randomly and then creates $K$ datasets with imputed values for incomplete variables. The order in which the variables are imputed is determined by the proportion of missing cases, starting with the variable that has the least missing data (van Buuren, 2018). Each variable uses the observed and most recently imputed values of the remaining ones, and this method is iterated on each dataset until convergence.
@@
After obtaining $K$ imputed datasets, the set of parameters of interest $\Psi$ is estimated separately on each of these datasets, resulting in a vector of estimates $(\hat{\Psi}_{1},...,\hat{\Psi}_{K})$ and estimated covariance matrices $(\hat{\Sigma}_{\Psi 1},...,\hat{\Sigma}_{\Psi K})$. Finally, the pooled estimate is obtained by combining the $K$ estimates using Rubin’s rule (Rubin, 2004)
@@
$\hat{\Sigma}_{\Psi}$ $=\frac{1}{K}\sum\limits_{k=1}^{K}\Sigma_{\Psi k}+$
$+\left(1+\frac{1}{K}\right)\frac{1}{K-1}\sum\limits_{k=1}^{K}(\Psi_{k}-\hat{\Psi})(\Psi_{k}-\hat{\Psi})^{T}$